# A Separable Self-attention Inspired by the State Space Model for Computer Vision

### v1

@@ III-D VMINet
IV Experiments
This section presents our experimental results, starting with the ImageNet classification task and then transferring the trained model to various downstream tasks, including object detection and instance segmentation.
IV-A Image Classification
Settings. We train the models on ImageNet-1K and evaluate the performance on ImageNet-1K validation set. For fair comparisons, our training settings mainly follow Vim. Specifically, we apply random cropping, random horizontal flipping, label-smoothing regularization, mixup, and random erasing as data augmentations. When training on $224\times 224$ input images, we employ AdamW with a momentum of 0.9 and a weight decay of 0.025 to optimize models. During testing, we apply a center crop on the validation set to crop out $224\times 224$ images. We train the VMINet models for 300 epochs using a cosine schedule. Unlike Vim, our experiments are performed on 3 A6000 GPUs. Therefore, we adjusted the total batch size and the initial learning rate to 384 and $5\times 10^{-4}$ respectively.
Results. We selected advanced CNNs, ViTs, and ViMs with comparable parameters and computational costs in recent years to compare with our method, and the results are shown in Table II. The various variants of VMINet are identical in every aspect except for the difference in embedding width. The experimental results demonstrate that VMINet overwhelmingly outperforms Vim [1], which utilizes a pure Mamba encoder. Compared to PlainMamba [18], which also employs depthwise convolutions, our VMINet exhibits significant advantages in terms of performance, efficiency, and model complexity. This suggests that VMI-SA is more suitable for visual tasks than Mamba. Furthermore, although VMINet is a demonstrative network architecture that has not been meticulously designed, it still achieves competitive results across various scales, particularly in lightweight scenarios. This suggests that the analysis presented in the previous sections may serve as a guiding principle, potentially reducing the unnecessary attempts researchers might make when designing attention mechanisms or general visual backbone networks.
TABLE III: Object detection and instance segmentation results on COCO.
[…]
ResNet-18 31M 207G 34.0 54.0 36.7 31.2 51.0 32.7
Vim-Ti 27M 189G 36.6 59.4 39.2 34.9 56.7 37.3
PVT-T 33M 208G 36.7 59.2 39.3 35.1 56.7 37.3
ResNet-50 44M 260G 38.0 58.8 41.4 34.7 55.7 37.2
@@ IV-A Image Classification
EfficientVMamba-S 31M 197G 39.3 61.8 42.6 36.7 58.9 39.2
ResNet-101 63M 336G 40.0 60.5 44.0 36.1 57.5 38.6
Vim-S 44M 272G 40.9 63.9 45.1 37.9 60.8 40.7
Swin-T 48M 267G 42.7 65.2 46.8 39.3 62.2 42.2
@@
Swin-S 69M 354G 44.8 66.6 48.9 40.9 63.2 44.2
IV-B Ablation Study
Setting aside the design philosophy, due to structural similarities, a reasonable suspicion is that the superior performance of VMINet may primarily be attributed to the introduction of depthwise separable convolutions. As shown in Fig. 3, for VMINet-S, after removing attention-related operations such as element-wise matrix multiplication and context vector generation, VMI-SA degenerates into a block similar to a ConvNeXt block [23]. Although this slightly reduces the number of parameters and computational complexity, the accuracy decreases from 80.2% to 77.6%.
IV-C Object Detection and Instance Segmentation
@@
Results. We summarize the comparison results of VMINet with other backbones in Table III. It can be seen that our VMINet consistently outperforms Vim. Similar to the results on classification tasks, VMINet achieves a good balance between the number of parameters and computational cost, achieving comparable results with advanced CNNs and ViTs.
V Conclusion
This paper presents a separable self-attention inspired by the visual Mamba (VMI-SA), which has a linear complexity. Through analysis and derivation, we believe that the element-wise multiplication operation used by separable self-attention can map the original features to a high-dimensional space for processing, which is more efficient than the matrix multiplication operation used by the classical softmax self-attention. Inspired by the Mamba design philosophy, we first establish local relevance through depthwise convolution, then limit the receptive field to the previous token, and then integrate local and global information according to the recursive state-space model, finally restoring the global receptive field and presenting the matrix form of VMI-SA. We believe that our work can provide a new perspective for the design of future attention mechanisms, that is, by changing the expression and constraints under a unified theoretical framework, to integrate the advantages of different methods. Currently, the research on VMI-SA is still in its infancy, and we believe that with reasonable network structure design, VMI-SA can further improve performance. In addition, the recursive form of VMI-SA is suitable for processing causal data and may be able to compete with other advanced methods in other fields.

### v2

@@ 4.2 Empirical studies on ImageNet-1K
ResNet-18 [6] 31M 207G 34.0 54.0 36.7 31.2 51.0 32.7
Vim-Ti [28] 27M 189G 36.6 59.4 39.2 34.9 56.7 37.3
PVT-T [23] 33M 208G 36.7 59.2 39.3 35.1 56.7 37.3
ResNet-50 [6] 44M 260G 38.0 58.8 41.4 34.7 55.7 37.2
@@
EfficientVMamba-S [17] 31M 197G 39.3 61.8 42.6 36.7 58.9 39.2
ResNet-101 [6] 63M 336G 40.0 60.5 44.0 36.1 57.5 38.6
Vim-S [28] 44M 272G 40.9 63.9 45.1 37.9 60.8 40.7
Swin-T [12] 48M 267G 42.7 65.2 46.8 39.3 62.2 42.2
@@
We use VMINet-XS without the mask matrix as the baseline model and apply different types of mask matrices to it separately. As shown in Table 4, even when using a highly sparse block diagonal matrix as the mask matrix, the model performance still shows a significant improvement. Experimentally, there is no difference in performance when using a banded matrix or a lower triangular matrix as the mask matrix. In addition, we explore the hybrid use of mask matrices. Specifically, the VMI-SA blocks in Stage 1 and Stage 2 use a banded matrix as the mask matrix, while the VMI-SA blocks in Stage 3 and Stage 4 alternately use lower triangular and banded matrices as the mask matrices. The experimental results demonstrate that the hybrid use of different types of mask matrices can achieve better performance. We speculate that carefully designed mask matrices can further enhance the performance of VMINet, and both structural design and parameterization are promising research directions.
4.3 Object Detection and Instance Segmentation on COCO
@@
Results. We summarize the comparison results of VMINet with other backbones in Table 5. It can be seen that our VMINet consistently outperforms Vim. Similar to the results on classification tasks, VMINet achieves a good balance between the number of parameters and computational cost, achieving comparable results with advanced CNNs and ViTs.
4.4 Semantic Segmentation on ADE20K
Settings. Following Vim [28], we train UperNet [25] with our VMINet on ADE20K dataset. In training, we employ AdamW with a weight decay of 0.01, and a total batch size of 16 to optimize models. The employed training schedule uses an initial learning rate of $6\times 10^{-5}$, linear learning rate decay, a linear warmup of 1500 iterations, and a total training of 160K iterations.
Table 6: Results of semantic segmentation on ADE20K.
Backbone Params mIoU
ResNet-50 [6] 67M 40.7
Vim-Ti [28] 34M 41.0
VMINet-XS (ours) 34M 42.7
Vim-S [28] 57M 44.1
Swin-T [12] 60M 44.4
VMINet-S (ours) 47M 44.8
ConvNeXt-T [13] 60M 46.7
VMINet-B (ours) 53M 47.2
Results. The results are presented in Table 6. Compared with Vim [28], VMINet once again demonstrates higher accuracy and outperforms models such as ResNet [6], Swin[12], and ConvNeXt [13], further validating the effectiveness of VMI-SA.
5 Conclusion
This paper presents a separable self-attention inspired by the visual Mamba (VMI-SA), with linear complexity. Through analysis and derivation, we demonstrate that the element-wise multiplication operation used by separable self-attention can also map the original features to a high-dimensional space for processing, which is more efficient than the matrix multiplication operation used by the classical softmax self-attention. Inspired by the Mamba design philosophy, we first establish local relevance through depthwise convolution, then limit the receptive field to the previous token, and then integrate local and global information according to the recursive state-space model to derive the recurrent form of VMI-SA. Considering the efficiency of matrix operations, we restore the global receptive field and present the matrix form of VMI-SA. We believe that our work can provide a new perspective for the design of future attention mechanisms, that is, by changing the expression and constraints under a unified theoretical framework, to integrate the advantages of different methods. Currently, the research on VMI-SA is still in its infancy, and we believe that with reasonable network structure design, VMI-SA can further improve performance. In addition, the recursive form of VMI-SA is suitable for processing causal data and may be able to compete with other advanced methods in other fields.